[figure: 9.7 cm]

THE INFLUENCE OF SUBSTITUTIONS ON THE MAGNETIZATION PROCESSES IN COMPOUNDS BASED ON

DyMn_{6}Ge_{6}

A.E.Bogdanov^{a,b}, A.V.Morozkin^{a}, S.A.Nikitin^{a,b}, T.I.Ivanova^{a,b},

J.K.Warchulska^{b} a Moscow State University, 119992 Moscow, Russia b International Laboratory of High Magnetic Fields and Low Temperatures, Gajowicka 95, 53-421 Wroclaw, Poland

In order to study the influence of substitutions in the rare-earth, manganese and germanium sublattices on the magnetic properties of DyMn_{6}Ge_{6} compound new solid solutions Dy_{0.8}Sm_{0.2}Mn_{6}Ge_{6}, DyMn_{5}CoGe_{6}, DyMn_{5}FeGe_{6} and DyMn_{6}Ge_{5}Si have been prepared. The magnetization isotherms of their polycrystalline samples were obtained in the temperature range 4.2-300 K and in external magnetic fields up to 140 kOe. The common feature of all isotherms is that no saturation has been reached up to the maximum value of the magnetic field. The magnetization curves for samples containing Co and Fe exhibit hysteresis loops below 20 K that are not observed for the initial DyMn_{6}Ge_{6} sample. It should be noted that some of the obtained isotherms of samples containing Co, Fe and Si are of almost the same linear character. This fact is probably connected with the predominance of ferrimagnetic or antiferromagnetic ordering with a weak ferromagnetic component. The higher values of magnetization of Dy_{0.8}Sm_{0.2}Mn_{6}Ge_{6} in comparison with non-substituted compound are due to antiparallel orientation of Sm and Dy magnetic moments. This research has been supported by RFBR grant N 00-02-17862.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : A.E.Bogdanov

Address for correspondence : Physics Department, Moscow State University, 119992 Moscow, Russia.

Email address : bogdanov@rem.phys.msu.su